OpenStudy (anonymous):

simplify the following expression square root of -104 and the square root of -8

OpenStudy (anonymous):

\[\sqrt{-104}=i2\sqrt{26}\]

OpenStudy (anonymous):

\[\sqrt{-8} = i2\sqrt2\]

OpenStudy (anonymous):

little confused here

OpenStudy (anonymous):

A negative inside a sqrt comes out as i...always

OpenStudy (anonymous):

104=2 x 2 x 2 x 13

OpenStudy (anonymous):

There is a single pair of 2s, so they come out, leaving you with 2 X 13 on the inside

OpenStudy (anonymous):

That gives you \[2i\sqrt{26}\]for the first one

OpenStudy (anonymous):

Same process applies to the second one.